# Security of Language Models for Code: A Systematic Literature Review

### v1

@@ 4.2.4. Summary of Existing Backdoor Attacks
Anand et al. (Anand et al., 2021a) define five classes of adversarial examples, including variable change, redundancy removal, synonym replacement, voice conversion, and variable interchange. Simultaneously, they conduct adversarial attacks on ALGOLISP generative models and Transformer, to uncover the limitations of code generation model robustness. They utilize Levenshtein distance, the ratio of Levenshtein distance to sentence length, and BERT-based distance metrics to measure the degree of perturbation in adversarial attacks. Experimental results demonstrate that, compared to traditional neural code generation models, the Transformer architecture exhibits stronger robustness. They also find that the original ALGOLISP dataset may introduce biases during the generation process.
Yang et al. (Yang et al., 2022) introduce ALERT, a black-box adversarial example attack that can efficiently transform inputs to cause erroneous outputs in victim models. ALERT, similar to MHM (Zhang et al., 2020a), modifies code variables to generate adversarial examples. Compared to MHM, ALERT pays more attention to the naturalness requirements of adversarial examples. Specifically, ALERT first uses CodeBERT and GraphCodeBERT pre-trained models to generate candidate alternative words for code variables. It then ranks these candidate alternative words based on cosine similarity and uses a greedy algorithm or a genetic algorithm to select adversarial examples from the candidate alternative words. Therefore, ALERT generates adversarial examples that retain the natural semantics and operational semantics of the code. Experiments demonstrate that ALERT outperforms MHM in attacking CodeBERT and GraphCodeBERT in vulnerability prediction, clone detection, and source code authorship attribution tasks.
@@
Li et al. (Li et al., 2022a) propose a robust defense method against known and unknown adversarial attacks to enhance the security of CodeLMs. They introduce two black-box adversarial attack methods specifically targeting the attribution of source code authors. These methods are Targeted Autoencoding Style Mimicry Attack and Untargeted Autoencoding Style Concealment Attack. The newly proposed adversarial attack methods leverage systematic preservation of semantic code style attributes and transformations in both attacks. Targeted Autoencoding Style Mimicry Attack: This attack extracts code style attributes from all code of the target author, synthesizes a comprehensive code style attribute, and identifies the code style attributes of the attacker’s code. Finally, it transforms the attacker’s code to mimic the style of the target code. Untargeted Autoencoding Style Concealment Attack: This attack extracts code style attributes from the attacker’s code, obtains code style attributes from other authors, identifies the differential set between the attacker’s code style attributes and those of other authors, selects the author with the highest probability of misattribution, and transforms the attacker’s code style to that of the selected target author. Experiments reveal that existing attribution models based on CodeLMs are highly vulnerable to known and unknown adversarial attacks, with untargeted attacks exhibiting significantly higher success rates than targeted attacks.
@@
Choi et al. (Choi et al., 2022) propose an efficient black-box adversarial attack method, called TABS, to reduce the search space. Specifically, TABS adopts beam search to find better adversarial examples and contextual semantic filtering to effectively reduce the search space. Experiments demonstrate that TABS can successfully launch adversarial attacks on the code search task of CodeBERT. Compared to Textfooler (Jin et al., 2020), BERT-Attack (Li et al., 2020a), and CLARE (Li et al., 2020b), TABS shows good performance in terms of attack success rate, number of queries, and other aspects.
@@
Jha et al. (Jha and Reddy, 2023) introduce a black-box adversarial attack method named CodeAttack, aiming to generate efficient but inconspicuous adversarial code examples using code structure. CodeAttack identifies the most vulnerable markers in the code and employs greedy search to generate alternative markers, replacing vulnerable markers with alternatives that minimally perturb the sample, maintain code consistency and fluency, and adhere to code constraints. Experimental results demonstrate that compared to TextFooler (Jin et al., 2020) and BERT-Attack (Li et al., 2020a), CodeAttack (Jha and Reddy, 2023) efficiently launches adversarial attacks on CodeT5, CodeBRER, and GraphCodeBERT in code transformation, code repair, and code summarization tasks, producing adversarial examples with higher syntactic correctness.
Gao et al. (Gao et al., 2023) propose a novel adversarial attack named DaK, specifically designed for CodeLMs. DaK creates discrete adversarial examples by preserving semantic code input through program transformations. DaK comprises three key components: Destroyer, Finder, and Merger. Destroyer alters the functionality of input code to weaken the features on which the model relies for predictions, using only semantically preserving transformations. Finder identifies a set of code predicted as the target label, computes critical features within this set, and selects features deemed the strongest and most concentrated in influencing the model’s prediction. Finally, Merger injects the selected critical features into the target program, generating discrete adversarial examples. Experiments on Code2Vec, GGNN, and CodeBERT demonstrate that DaK outperforms Imitator (Quiring et al., 2019) and DAMP (Yefet et al., 2020) on attacking undefended or unreinforced CodeLMs.
Nguyen et al. (Nguyen et al., 2023) propose a novel black-box adversarial attack framework called GraphCodeAttack. To assess the robustness of CodeLMs, GraphCodeAttack automatically discovers important code patterns influencing model decisions to perturb the code structure of input models. Specifically, GraphCodeAttack identifies discriminative abstract syntax tree (AST) patterns that influence model decisions from source code and target model outputs. It then selects suitable patterns to insert dead code into the source code, creating adversarial examples. Experimental results show that GraphCodeAttack achieves an average attack success rate of 0.841 on author attribution, code vulnerability detection, and code clone detection tasks for CodeBERT and GraphCodeBERT CodeLMs, surpassing ALERT and CARROTA adversarial attack methods.
Yu et al. (Yu et al., 2023) introduce an efficient black-box adversarial attack framework, ADVULCODE, specifically designed for generating code adversarial examples in the context of vulnerability detection tasks based on CodeLMs. Unlike ALERT (Yang et al., 2022) and MHM (Zhang et al., 2020a), which generate adversarial examples by perturbing identifiers, ADVULCODE utilizes equivalent transformation rules that do not alter the semantic meaning of the source code. It targets susceptible code segments for transformation, thus reducing the overall perturbation level of the source code. Specifically, ADVULCODE employs equivalent transformations to generate candidate code statements and introduces an improved Monte Carlo search tree to guide the selection of candidate statements and generate adversarial examples. In comparison to ALERT, ADVULCODE can handle models with identifier name mappings. Experimental results demonstrate that ADVULCODE outperforms ALERT in adversarial attacks for vulnerability detection tasks.
@@
Na et al. (Na et al., 2023) propose an efficient and advanced black-box adversarial attack method named DIP, which does not require any information about the pre-trained deep models and avoids additional training computations. DIP creates adversarial examples by inserting dead code into the source code. Initially, DIP leverages the pre-trained CodeBERT model to identify vulnerable positions in the source code. It then utilizes cosine similarity to find candidate code snippets in the code repository that are least semantically similar to the source code. Finally, DIP selects the code statement with minimal disturbance from these candidate code snippets, wraps it with an unused variable, and inserts it as dead code into the vulnerable position of the source code. Since DIP launches adversarial attacks by inserting dead code, it preserves the functionality and compilability of the source code. Experiments show that DIP achieves the best CodeBLEU (Ren et al., 2020) on nine target models and the best disturbance ratio on most target models.
Zhang et al. (Zhang et al., 2023c) propose RNNS, a black-box adversarial attack method designed to assess the robustness of pre-trained CodeLMs by generating adversarial test data based on model uncertainty and output variation. Similar to MHM (Zhang et al., 2020a) and ALERT (Yang et al., 2022), RNNS generates adversarial examples by renaming local variables in the source code. In contrast to MHM and ALERT, RNNS collects variable names from real code to generate statistically characteristic adversarial examples while pursuing a high attack success rate. Specifically, RNNS first constructs a search space for variable names based on semantic similarity using real code data. Then, it utilizes a Representation Nearest Neighbor Search method to find adversarial examples from the namespace guided by model output changes. Failed attack samples guide the next round of attacks in RNNS. Experimental results indicate that, compared to MHM and ALERT, RNNS generates adversarial instances with smaller perturbations and achieves superior attack effectiveness on CodeBERT, GraphCodeBERT, and CodeT5 across six code tasks in three programming languages.
@@
Abuhamad et al. (Abuhamad et al., 2023) introduce SHIELD to attack the code authorship attribution system, demonstrating the insufficient robustness of current code authorship attribution systems. SHIELD injects carefully selected code samples into the target code and then applies Stunnix  [^1] http://stunnix.com for code-to-code obfuscation. SHIELD employs obfuscation to make it challenging for models to statically identify or remove the injected code portions, rather than solely conceal the identity of the code author. Adversarial samples are inputted into a black-box code authorship attribution system, and then the probability distribution of the identification model is modified by SHIELD through altering the code expressions and statements of the adversarial samples. SHIELD achieves a success rate of over 95.5% and 66% in non-targeted and targeted attacks, respectively, with a decrease in model identification confidence of over 13%. This indicates that the current code authorship attribution system is highly susceptible to the influence of adversarial samples.
@@
Zhang et al. (Zhang et al., 2024c) propose a novel black-box adversarial attack technique to address the inefficiency of adversarial attacks based on simple search strategies, named CodeBERT-attack. CodeBERT-attack iteratively performs identifier replacement attacks. Firstly, CodeBERT-attack locates vulnerable identifiers in the code through masked classification. Then, it utilizes CodeBERT to generate new perturbed identifier names, introducing variations to the identifiers while preserving the overall code structure. Finally, it tests the perturbed samples on the victim model, selecting the samples with the lowest change in prediction accuracy on the ground true labels. Their experiments demonstrate that CodeBERT-attack can reduce the accuracy of LSTM and CodeBERT by an average of 30% in function classification and code clone detection tasks. Moreover, the paper indicates that pre-training can increase the model’s adversarial robustness.
@@
Yang et al. (Yang et al., 2024e) propose RADAR-Attack, a gradient-free adversarial attack method designed to expose the vulnerabilities of CodeLMs by manipulating method names. RADAR-Attack targets CodeLMs with adversarial method names that are visually and semantically similar to the original inputs but cause the model to generate incorrect code. RADAR-Attack leverages a two-step approach: first, it generates candidate method names using semantic and visual similarity techniques, and second, it applies a genetic algorithm to optimize these candidates, minimizing the performance of the model as measured by CodeBLEU. This gradient-free approach makes RADAR-Attack highly effective even in black-box scenarios where the internal details of the model are unknown. Experiments demonstrate that RADAR-Attack can significantly reduce performance in models like CodeGPT, PLBART, and CodeT5 in fine-tuning tasks, and in Replit, CodeGen, and CodeT5+ in zero-shot tasks, showcasing its capability to launch effective adversarial attacks against CodeLMs.
@@
Awal et al. (Awal et al., 2024) propose a novel adversarial attack method aimed at assessing the robustness of LM in software analytics tasks. The approach minimizes perturbations in feature space, leveraging ML explainability techniques to identify the most influential features impacting model decisions. It emphasizes that adversarial example generation in software analytics fundamentally differs from methods in computer vision, as any modifications to feature values are easily perceptible. The study employs the reverse elbow method to determine optimal feature combinations. The reverse elbow method evaluates the point at which the improvement in performance metric values ensures that generated adversarial examples induce misclassifications while maintaining feature validity. Experimental results demonstrate that the generated adversarial examples effectively reduce model accuracy, validating the method’s potential for enhancing model security.
Wu et al. (Wu et al., 2023) propose a novel attack method, DeceptPrompt, targeting LLM-driven code generation models. DeceptPrompt generates adversarial natural language instructions that drive these models to produce functionally correct but vulnerable code. Specifically, given a benign prompt and a target vulnerability, DeceptPrompt introduces an evolution-based optimization framework to modify the prompt with a semantic-preserving prefix or suffix. The framework systematically manipulates the input to induce the model into generating vulnerable code, while maintaining the correct functionality. Extensive experiments demonstrate that DeceptPrompt can attack multiple popular LLMs, such as CodeLlama, StarCoder, and WizardCoder, achieving a success rate of up to 62.5% in targeted vulnerability injection across C and Python programming languages. These results reveal significant security weaknesses in current LLM-driven code generation models, with DeceptPrompt successfully bypassing security checks and generating vulnerable code in realistic deployment scenarios.
Zhang et al. (Zhang et al., 2024d) propose a novel attack method, STRUCK, targeting CodeLMs for code representation. STRUCK fully exploits the structural features of code to generate adversarial examples that mislead CodeLMs while maintaining the original semantics of the code. STRUCK integrates both global and local perturbation methods, using hierarchical structural features to select appropriate perturbations during adversarial example generation. Through comprehensive experiments on several code classification tasks, STRUCK demonstrates its superior effectiveness, efficiency, and imperceptibility compared to existing attacks like MHM and S-CARROTA (Zhang et al., 2022a). The experimental results show that STRUCK achieves a performance degradation of over 80% on various CodeLMs and significantly improves the robustness of models when used for adversarial training.
@@
Yang et al. (Yang et al., 2024a) propose CodeTAE, an adversarial attack method targeting transfer learning in source code models. CodeTAE crafts adversarial examples that exploit the transferability of pre-trained code encoders by introducing semantic-preserving transformations. Unlike previous methods, CodeTAE operates in a cross-domain setting, where the attacker only has access to an open-source pre-trained encoder without any information about the victim model’s downstream task or training data. Using a genetic algorithm, CodeTAE generates substitute identifiers and applies obfuscations to induce misclassifications in source code models, achieving attack success rates ranging from 30% to 80%. Experiments on multiple code classification tasks, such as authorship attribution and clone detection, demonstrate the attack’s effectiveness, as well as its utility for adversarial fine-tuning to enhance model robustness.
@@
Yao et al. (Yao et al., 2024) propose a black-box adversarial attack method, CARL, which leverages reinforcement learning to generate adversarial examples for code-based pre-trained language models (PLMs). CARL consists of a programming language encoder and a perturbation prediction layer, which are initialized with a pre-trained masked CodeBERT model. The method formulates the task of generating adversarial examples as a sequence decision-making problem and optimizes it through policy gradient techniques. CARL uses a combination of reward functions, including attack effectiveness and attack quality, to guide the learning process. Unlike traditional search-based approaches, CARL performs a one-step perturbation prediction, making it significantly faster during inference. Extensive experiments on tasks like code summarization, code translation, and code refinement show that CARL achieves higher success rates while maintaining code quality.
@@
Zhou et al. (Zhou et al., 2024) propose a multi-objective adversarial attack method (MOAA), which uses evolutionary multi-objective optimization to generate adversarial examples for code representation models. Unlike previous methods, MOAA integrates the NSGA-II algorithm with CodeT5 to generate adversarial examples that preserve both syntactic and semantic integrity. MOAA targets DNN models used for tasks such as code classification, clone detection, and code summarization, ensuring high-quality adversarial examples with minimal perturbation. Extensive experiments across five tasks and ten datasets reveal that MOAA produces diverse and transferable adversarial examples, with successful attacks on models like CodeBERT, GraphCodeBERT, and CodeT5. The results highlight MOAA’s effectiveness in generating robust adversarial examples while offering insights into the vulnerabilities of pre-trained code models.
@@
He et al. (He et al., 2024) propose MalwareTotal, a novel adversarial attack method targeting static malware detection models. MalwareTotal generates adversarial malware variants by leveraging sequence-aware bypass tactics that preserve the functionality of the malware while applying multi-faceted manipulations. The method employs proximal policy optimization to identify vulnerable manipulation sequences, effectively bypassing multiple detection mechanisms, including ML-based, signature-based, and hybrid anti-malware software.
4.3.4. Summary of Existing Adversarial Attacks

### v2

@@ 4.3.3. Black-box Attacks on CodeLMs
To reveal the limitations of the robustness of code generation models, Anand et al. (Anand et al., 2021a) define five classes of adversarial examples, including variable change, redundancy removal, synonym replacement, voice conversion, and variable interchange, and conduct adversarial attacks on ALGOLISP generative models and Transformer. They utilize Levenshtein distance, the ratio of Levenshtein distance to sentence length, and BERT-based distance metrics to measure the degree of perturbation in adversarial attacks. Experimental results demonstrate that, compared to traditional neural code generation models, the Transformer architecture exhibits stronger robustness. They also find that the original ALGOLISP dataset may introduce biases during the generation process.
To further investigate the security of LLM-driven code generation models, Wu et al. (Wu et al., 2023) propose a attack method, DeceptPrompt. DeceptPrompt generates adversarial natural language instructions that drive these models to produce functionally correct but vulnerable code. Specifically, given a benign prompt and a target vulnerability, DeceptPrompt introduces an evolution-based optimization framework to modify the prompt with a semantic-preserving prefix or suffix. The framework systematically manipulates the input to induce the model into generating vulnerable code, while maintaining the correct functionality. Extensive experiments demonstrate that DeceptPrompt can attack multiple popular LLMs, such as CodeLlama, StarCoder, and WizardCoder, achieving a success rate of up to 62.5% in targeted vulnerability injection across C and Python programming languages. These results reveal significant security weaknesses in current LLM-driven code generation models, with DeceptPrompt successfully bypassing security checks and generating vulnerable code in realistic deployment scenarios.
@@
Li et al. (Li et al., 2022a) introduce two black-box adversarial attack methods specifically targeting the attribution of source code authors. These methods are Targeted Autoencoding Style Mimicry Attack and Untargeted Autoencoding Style Concealment Attack. The proposed adversarial attack methods leverage systematic preservation of semantic code style attributes and transformations in both attacks. Targeted Autoencoding Style Mimicry Attack: This attack extracts code style attributes from all code of the target author, synthesizes a comprehensive code style attribute, and identifies the code style attributes of the attacker’s code. Finally, it transforms the attacker’s code to mimic the style of the target code. Untargeted Autoencoding Style Concealment Attack: This attack extracts code style attributes from the attacker’s code, obtains code style attributes from other authors, identifies the differential set between the attacker’s code style attributes and those of other authors, selects the author with the highest probability of misattribution, and transforms the attacker’s code style to that of the selected target author. Experiments reveal that existing attribution models based on CodeLMs are highly vulnerable to known and unknown adversarial attacks, with untargeted attacks exhibiting significantly higher success rates than targeted attacks.
@@
To reduce the search space for generating adversarial examples, Choi et al. (Choi et al., 2022) propose TABS. Specifically, TABS adopts beam search to find better adversarial examples and contextual semantic filtering to reduce the search space. Experiments demonstrate that TABS can successfully launch adversarial attacks on the code search task of CodeBERT. Compared to Textfooler (Jin et al., 2020), BERT-Attack (Li et al., 2020a), and CLARE (Li et al., 2020b), TABS shows good performance in terms of attack success rate, number of queries, and other aspects.
@@
Jha et al. (Jha and Reddy, 2023) introduce a black-box adversarial attack method named CodeAttack, aiming to generate inconspicuous adversarial code examples using code structure. CodeAttack identifies the most vulnerable markers in the code and employs greedy search to generate alternative markers, replacing vulnerable markers with alternatives that minimally perturb the sample, maintain code consistency and fluency, and adhere to code constraints. Experimental results demonstrate that compared to TextFooler (Jin et al., 2020) and BERT-Attack (Li et al., 2020a), CodeAttack (Jha and Reddy, 2023) efficiently launches adversarial attacks on CodeT5, CodeBRER, and GraphCodeBERT in code transformation, code repair, and code summarization tasks, producing adversarial examples with higher syntactic correctness.
@@
Na et al. (Na et al., 2023) propose a black-box adversarial attack method named DIP, which does not require any information about the pre-trained deep models and avoids additional training computations. DIP creates adversarial examples by inserting dead code into the source code. Initially, DIP leverages the pre-trained CodeBERT model to identify vulnerable positions in the source code. It then utilizes cosine similarity to find candidate code snippets in the code repository that are least semantically similar to the source code. Finally, DIP selects the code statement with minimal disturbance from these candidate code snippets, wraps it with an unused variable, and inserts it as dead code into the vulnerable position of the source code. Since DIP launches adversarial attacks by inserting dead code, it preserves the functionality and compilability of the source code. Experiments show that DIP achieves the best CodeBLEU (Ren et al., 2020) on nine target models and the best disturbance ratio on most target models.
@@
To reveal the insufficient robustness of current code authorship attribution systems, Abuhamad et al. (Abuhamad et al., 2023) introduce SHIELD. SHIELD injects carefully selected code samples into the target code and then applies Stunnix (Stunnix, 2002) for code-to-code obfuscation. SHIELD employs obfuscation to make it challenging for models to statically identify or remove the injected code portions, rather than solely conceal the identity of the code author. Adversarial samples are inputted into a black-box code authorship attribution system, and then the probability distribution of the identification model is modified by SHIELD through altering the code expressions and statements of the adversarial samples. SHIELD achieves a success rate of over 95.5% and 66% in non-targeted and targeted attacks, respectively, with a decrease in model identification confidence of over 13%.
@@
To address the inefficiency of adversarial attacks based on simple search strategies, Zhang et al. (Zhang et al., 2024c) propose CodeBERT-attack, a black-box adversarial attack technique. CodeBERT-attack iteratively performs identifier replacement attacks. Firstly, CodeBERT-attack locates vulnerable identifiers in the code through masked classification. Then, it utilizes CodeBERT to generate new perturbed identifier names, introducing variations to the identifiers while preserving the overall code structure. Finally, it tests the perturbed samples on the victim model, selecting the samples with the lowest change in prediction accuracy on the ground true labels. Their experiments demonstrate that CodeBERT-attack can reduce the accuracy of LSTM and CodeBERT by an average of 30% in function classification and code clone detection tasks. Moreover, the paper indicates that pre-training can increase the model’s adversarial robustness.
@@
Yang et al. (Yang et al., 2024g) propose RADAR-Attack, a gradient-free adversarial attack method designed to expose the vulnerabilities of CodeLMs by manipulating method names. RADAR-Attack targets CodeLMs with adversarial method names that are visually and semantically similar to the original inputs but cause the model to generate incorrect code. RADAR-Attack leverages a two-step approach: first, it generates candidate method names using semantic and visual similarity techniques, and second, it applies a genetic algorithm to optimize these candidates, minimizing the performance of the model as measured by CodeBLEU. This gradient-free approach makes RADAR-Attack highly effective even in black-box scenarios where the internal details of the model are unknown. Experiments demonstrate that RADAR-Attack can significantly reduce performance in models like CodeGPT, PLBART, and CodeT5 in fine-tuning tasks, and in Replit, CodeGen, and CodeT5+ in zero-shot tasks, showcasing its capability to launch effective adversarial attacks against CodeLMs.
@@
Awal et al. (Awal et al., 2024) propose a adversarial attack method aimed at assessing the robustness of LM in software analytics tasks. The approach minimizes perturbations in feature space, leveraging ML explainability techniques to identify the most influential features impacting model decisions. It emphasizes that adversarial example generation in software analytics fundamentally differs from methods in computer vision, as any modifications to feature values are easily perceptible. The study employs the reverse elbow method to determine optimal feature combinations. The reverse elbow method evaluates the point at which the improvement in performance metric values ensures that generated adversarial examples induce misclassifications while maintaining feature validity. Experimental results demonstrate that the generated adversarial examples effectively reduce model accuracy, validating the method’s potential for enhancing model security.
@@
Yang et al. (Yang et al., 2024a) propose CodeTAE, an adversarial attack method targeting transfer learning in source code models. CodeTAE crafts adversarial examples that exploit the transferability of pre-trained code encoders by introducing semantic-preserving transformations. CodeTAE operates in a cross-domain setting, where the attacker only has access to an open-source pre-trained encoder without any information about the victim model’s downstream task or training data. Using a genetic algorithm, CodeTAE generates substitute identifiers and applies obfuscations to induce misclassifications in source code models, achieving attack success rates ranging from 30% to 80%. Experiments on multiple code classification tasks, such as authorship attribution and clone detection, demonstrate the attack’s effectiveness, as well as its utility for adversarial fine-tuning to enhance model robustness.
@@
Yao et al. (Yao et al., 2025) propose a black-box adversarial attack method, CARL, which leverages reinforcement learning to generate adversarial examples for code-based pre-trained language models. CARL consists of a programming language encoder and a perturbation prediction layer, which are initialized with a pre-trained masked CodeBERT model. The method formulates the task of generating adversarial examples as a sequence decision-making problem and optimizes it through policy gradient techniques. CARL uses a combination of reward functions, including attack effectiveness and attack quality, to guide the learning process. Unlike traditional search-based approaches, CARL performs a one-step perturbation prediction, making it significantly faster during inference. Extensive experiments on tasks like code summarization, code translation, and code refinement show that CARL achieves higher success rates while maintaining code quality.
@@
Zhou et al. (Zhou et al., 2024) propose a multi-objective adversarial attack method, MOAA, which uses evolutionary multi-objective optimization to generate adversarial examples for code representation models. Unlike previous methods, MOAA integrates the NSGA-II algorithm with CodeT5 to generate adversarial examples that preserve both syntactic and semantic integrity. MOAA targets DNN models used for tasks such as code classification, clone detection, and code summarization, ensuring high-quality adversarial examples with minimal perturbation. Extensive experiments across five tasks and ten datasets reveal that MOAA produces diverse and transferable adversarial examples, with successful attacks on models like CodeBERT, GraphCodeBERT, and CodeT5. The results highlight MOAA’s effectiveness in generating robust adversarial examples while offering insights into the vulnerabilities of pre-trained code models.
@@
He et al. (He et al., 2024) propose MalwareTotal, an adversarial attack method targeting static malware detection models. MalwareTotal generates adversarial malware variants by leveraging sequence-aware bypass tactics that preserve the functionality of the malware while applying multi-faceted manipulations. The method employs proximal policy optimization to identify vulnerable manipulation sequences, effectively bypassing multiple detection mechanisms, including ML-based, signature-based, and hybrid anti-malware software.
4.3.4. Summary of Adversarial Attacks